OpenStudy (anonymous):

Evaluate cos2π/3−arctan 1 these are the 8 options answers 1.3√−1/2 2.1−√3/2 3.π+2/4 4.1−π/2 5.−π+2/4 6.π−2/4 7.−3√+2/4 8.−3√+1/2

hartnn (hartnn):

tan pi/4 =1 so arctan 1 = pi/4 could u find cos 2pi/3 ? or need help with that also ?

OpenStudy (anonymous):

yes pls

hartnn (hartnn):

do you know unit circle ?

OpenStudy (anonymous):

just a little

hartnn (hartnn):

there you will find cos 2pi/3 = -0.5 = -1/2

hartnn (hartnn):

so, overall answer u get is -1/2-pi/4

OpenStudy (anonymous):

thx

OpenStudy (anonymous):

one more question

OpenStudy (anonymous):

giving a circle of radius 2 and center at the point (−1,2)? solve for the equation

hartnn (hartnn):

ok, the circle with radius r and center x1,y1 is \[(x-x_1)^2(y-y_1)^2=r^2\] just plug in values.

OpenStudy (anonymous):

thx so much